a rectangle has a perimeter of 60 centimeters and the base is 1 less than twice the height what is the area of the rectangle?
Answer by jorel1380(3719)

2h+2(2h-1)=60
2h+4h-2=60
6h=62
h=62/6
b=118/6
Area=203.222 cc's..